GlueX Start Counter Meeting, July 24, 2014

From GlueXWiki
Jump to: navigation, search

GlueX Start Counter Meeting
Thursday, July 24, 2014
10:00 am EDT
CEBAF Center, Room F326/327


  1. Announcements
  2. Review minutes from July 10th meeting
  3. Electronics Trip Report (Nick)
  4. Detector Construction Status (Eric)
  5. Schedule Review (all)
    1. Readiness Review Date: August 4[?]
    2. Readiness Review Committee and Charge
    3. Video tour at Readiness Review?




Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2014-3Q on the JLab CUE. This directory is accessible from the web at .



  • FIU: Werner Boeglin, Eric Pooser
  • JLab: Eugene Chudakov, Mark Ito (chair), Bruce Lenzer, Nick Sandoval, Tim Whitlatch, Beni Zihlmann

Review of Minutes from July 10th Meeting

We looked at the minutes. Mark needs to check on the status of an itemized invoice from FIU.

Electronics Trip Report

Nick was at FIU last week, Monday through Thursday to install the start counter on-detector electronics. Overall the trip went well and all goals were accomplished.

  • ST1/ST2/ST3 boards all delivered.
  • MPOD for bias and low voltage was delivered and fired up.
  • A cosmic trigger with external counters was set-up and data taking started.

A couple of minor issues:

  • The JLab laptop revoked Nick's admin privileges when it was booted at FIU. A work-around was found and the control software was installed on two personal laptops (Nick's and Eric's).
  • The ST2 cables to the ST3 were too short to run out of the chicane of the dark box, so the ST3 enclosure had to be placed inside the dark box. LEDs on the ST3 were disconnected for obvious reasons.

Detector Construction Status

Eric gave the report.

  • Eric wrote a progress report and issued it as a GlueX Note. It describes the polishing procedure.
  • All 50 scintillators tested before polishing and characterized. The worst of them were used to develop the polishing procedure.
  • Polishing procedure finalized.
    • Polish one hour with 0.3 micron alumina suspension.
    • Test timing resolution.
    • Another hour at 0.05 micron if needed.
  • Can polish 2-3 a day, total should take 2-3 weeks to finish.
    • There are two polishing stations now.
  • 49 of the counters will be polished; one will be kept as-is for deterioration studies.
  • Several bars have been tested multiple times over the past two weeks; thus far there is no sign of deterioration.
  • Camera for adjusting alignment and separation of the scintillator and SiPMs have been calibrated.
  • Shimming will likely have to be done.
  • Werner ask about getting 25 micron Kapton from JLab to be used for shimming material.
  • The gap is set by using a spacer of foil and paper to set the position of the scintillator.

Light Tightening

Werner suggested that we do the final light tightening of the detector at JLab. All testing at FIU during construction can be done in the dark box at FIU. Beni pointed out that there is a large dark box in EEL 126 that can be used for testing the scheme here. Mark mentioned that Chuck had proposed the same idea in a conversation last week.

We decided to pursue this track. There are still several ideas around for the exact materials and procedure; they need to be finalized in advance of delivery of the detector to JLab.

Readiness Review Planning

The Review will be held on Monday, August 4, at 10 am. At JLab L207 is reserved.

The review committee consists of:

  • Beni
  • Bruce
  • Mark
  • Nick
  • Tim

We looked at the draft charge that Bruce circulated some time ago. Eric has been looking at that draft to guide his preparations. Bruce will be making changes to finalize the document.

Eric will lead us through the construction procedures at the review. [Added in press: it would be nice to get copies of the written procedures to the committee in advance of the review.]

Bruce had suggested some sort of video tour since there will not be any committee members on site at FIU. FIU will think about a plan for this; we thought that perhaps BlueJeans on a mobile device could be used.

Gain Issue

[Added in press: we neglected to discuss the reduction in gain of the pre-amps after cosmic tests at FIU last week. Reduction by about a factor of 5 is planned.]

Action Items

  1. Look into sending Kapton sheets to FIU for use in shimming. (Tim)
  2. Finalize the Review Charge document. (Bruce)
  3. Get an advance copy of the procedures of the committee. (Eric)
  4. Raise the gain issue at a future meeting. (Mark)